# Restricted Receptive Fields for Face Verification

### v1

@@ I Introduction
As opposed to post-hoc approaches, interpretable models emphasize constrained design choices to make decisions inherently more interpretable [13, 26, 27, 28, 29, 30]. One prominent example is ProtoPNet [31], which expresses model outputs as a weighted sum of similarities to learned prototypes derived from training images. This formulation enables localized and case-based reasoning by relating regions of the input image to the prototypes. However, its applicability is limited to closed-set recognition, as the prototypes are drawn from training data.
In this work, we propose a distance metric for measuring similarity between face images. Unlike the traditional approach of representing a face image with a single feature vector, we first extract representations from restricted receptive fields. Patch-level similarity scores are then computed between image pairs, and the global similarity score is defined as the sum of these patch similarities. This formulation provides a more human-understandable similarity assessment, as the overall score is decomposed into contributions from different sub-regions. Moreover, because patch similarities are an inherent part of the decision process, our method eliminates the need for post-hoc explanation techniques.
We evaluate our approach using two patch sizes, $28\times 28$ and $56\times 56$, on $112\times 112$ face images. Although restricted receptive fields constrain the available spatial context, we surprisingly find that this design choice improves verification accuracy compared to state-of-the-art methods when $56\times 56$ patches are used for similarity measurement. Notably, even with patches as small as $28\times 28$, our method achieves competitive verification performance.
This paper is organized as follows. In Section II, we review the differences between post-hoc explanation methods and inherently interpretable approaches. Section III introduces two approaches for measuring similarity from patch representations. First, in Section III-A, we define a global distance metric as a weighted sum of region-based similarities. Next, in Section III-B, we present RRFNet and show that with a small modification to the ResNet architecture, similarity decisions become more interpretable in terms of patch-level similarities. Experimental results are given in Section IV for two approaches. Finally, we summarize the proposed work and discuss future work in Section V.
@@ II Related Work
Refer to caption (0,0) Refer to caption (14,0) Refer to caption (28,0) Refer to caption (42,0) Refer to caption (56,0) Refer to caption (70,0) Refer to caption (84,0) Refer to caption (0,14) Refer to caption (14,14) Refer to caption (28,14) Refer to caption (42,14) Refer to caption (56,14) Refer to caption (70,14) Refer to caption (84,14) Refer to caption (0,28) Refer to caption (14,28) Refer to caption (28,28) Refer to caption (42,28) Refer to caption (56,28) Refer to caption (70,28) Refer to caption (84,28) Refer to caption (0,42) Refer to caption (14,42) Refer to caption (28,42) Refer to caption (42,42) Refer to caption (56,42) Refer to caption (70,42) Refer to caption (84,42) Refer to caption (0,56) Refer to caption (14,56) Refer to caption (28,56) Refer to caption (42,56) Refer to caption (56,56) Refer to caption (70,56) Refer to caption (84,56) Refer to caption (0,70) Refer to caption (14,70) Refer to caption (28,70) Refer to caption (42,70) Refer to caption (56,70) Refer to caption (70,70) Refer to caption (84,70) Refer to caption (0,84) Refer to caption (14,84) Refer to caption (28,84) Refer to caption (42,84) Refer to caption (56,84) Refer to caption (70,84) Refer to caption (84,84)    a) $28\times 28$ Refer to caption (0,0) Refer to caption (28,0) Refer to caption (56,0) Refer to caption (0,28) Refer to caption (28,28) Refer to caption (56,28) Refer to caption (0,56) Refer to caption (28,56) Refer to caption (56,56)   b) $56\times 56$ Refer to caption c) $112\times 112$
@@
Several components of interpretability in machine learning models, such as sparsity and linearity, were reviewed before the rise of feature learning approaches [45, 46, 47, 48]. Logistic regression is broadly employed for high-stakes decisions as it is considered interpretable. One of the important interpretable properties is the ability to express model decision globally, as the same coefficients are used for every decision. Linearity between output and input enables easy assessment of feature importance for humans. However, obtaining high level human interpretable features is a significant challenge for high dimensional input, such as images.
@@
A great amount of research has focused on post-hoc explanation approaches, where the decision of a black-box model is explained by generating heatmaps that highlight important pixels that influence the model’s predictions most. These visual explanations can help identify relevant regions, offering insights into where the model looks when making decisions. However, evaluating the quality of these explanations remains an unsolved challenge. Furthermore, post-hoc methods can be misleading, as similar heatmaps are often generated for both correct and incorrect predictions, undermining the reliability of these explanations [49, 50, 51, 52, 53, 9].
@@
Another line of research in recent years focuses on incorporating interpretability directly into the vision models [54, 31, 55, 56]. The main component of these models on visual recognition tasks is to decompose images into human-interpretable parts and make predictions based on a weighted combination of these parts. In [31], they employ a prototype learning approach to explain the decision through part-based evidence. Prototypes are assigned to each class based on representative image patches from the training set, and the final prediction is calculated as a weighted sum of similarity to these prototypes. While this part-based evidence framework enhances interpretability, it is restricted to classes present in the training set, limiting its generalization to unseen categories.
Recent works in face recognition have explored utilization of post-hoc explanation approaches [21, 22, 23, 24]. However, aforementioned limitations of post-hoc explanations remain. In contrast to post-hoc methods, our work aims to make its decisions inherently interpretable based on local similarities. Our approach allows straightforward evaluation of explanations, as error rates can be independently assessed for each local region. Furthermore, it does not require any additional post-hoc approaches as the explanations are integrated into the decision-process.
A similar idea is explored in BagNet [44], where the receptive field of a CNN is restricted by replacing standard $3\times 3$ kernels with $1\times 1$ kernels. This enables the network to learn local features, providing regional explanations of its decisions. However, replacing most of the filters with $1\times 1$ to control receptive field size results in significantly reduced number of parameters compared to modern CNN architectures, thereby improving interpretability at the expense of accuracy. While BagNet’s approach aligns with our work in terms of breaking down the decision process into restricted receptive fields, two key differences distinguish our method. First, BagNet focuses solely on object classification tasks with known classes from the training set, whereas our approach introduces an interpretable similarity metric for face verification that works with unseen subjects at test time. Second, while BagNet highlights a trade-off between interpretability and accuracy as their approach only considers using $1\times 1$ filters to control receptive field, our approach seeks to prioritize high accuracy without losing expressive power of CNNs. Our experiments demonstrate that simple design choices over the traditional approach can enhance model transparency while maintaining high verification performance.
@@ III Methodology
In this work, we propose a distance metric to decompose face similarity measure into local regions. For an image size of $W\times H$, we choose $w<W$ and $h<H$ to learn local representations for each $w\times h$ restricted patch. The image patches are uniformly distributed across the image. Two approaches are considered for learning local representations and constructing global image similarity metric: (i) training region-based CNNs for each $k$ patches (see Section III-A) and (ii) single CNN to learn a global image representation as the mean of local representations (see Section III-B). In the first approach, we adopt the same training objective as in [15] to learn local features for each patch obtained from $W\times H$ face images. After the CNNs are trained, a second training phase learns weights for each region similarity using logistic regression. Weighted sum of $k$ local similarity scores, obtained form corresponding regions between two images, is used to the build global distance metric. In the second approach, rather than training separate CNNs, a single CNN is trained and the global representation is computed as the mean of the patch-level features, enabling face similarity to be measured directly from patch representations. We use the terms patch and restricted receptive field interchangeably throughout the manuscript to describe our approach.
III-A Patch Representation Learning
We first divide each face image into a set of smaller, uniformly distributed patches. Each patch is treated as an independent unit, allowing the model to capture fine-grained localized feature representations. Unlike the traditional approach of generating a single $N$-dimensional representation for an entire image, we obtain $(k,N)$-dimensional representation where a face image is divided into $k$ patches. Patches are defined with position and window size. For a given face image, we extract $k$ patches, which we will refer to as $P_{1},P_{2},\dots,P_{k}$, each corresponding to a distinct region and size of the image. While $112\times 112$ is traditionally used image size for face recognition research in recent works, we analyze two patch sizes, $28\times 28$ and $56\times 56$ in our experiments, demonstrated in Figure 3.
@@ III-A Patch Representation Learning
Let the output of each CNN, trained in the first step, corresponding to the $i$-th patch be $f_{i}$, which is a $N$-dimensional feature vector representing the learned local features of patch $P_{i}$. Local similarities between two images is computed for each corresponding patches $P_{i}^{A}$ from image $A$ and $P_{i}^{B}$ image $B$ using cosine similarity:
@@ III-B Restricted Receptive Field Network (RRFNet)
Comparison of the proposed RRFNet approach with the ResNet [57] is depicted in Table I. While traditional approach with ResNet produce a $512$-dimensional representation from the entire face image, our approach extracts $512$-dimensional representations for each restricted receptive field. At the end of the RRFNet, a global representation is obtained as the mean of the patch representations enabling expression of the global similarity as patch similarities.
Patch-level Similarity Metric. Typically, cosine similarity between two $512$-dimensional face representations, extracted from a pretrained recognition model, is used for face verification. In our approach, the global face representation is defined as the mean of the patch-level representations. This formulation shows that the similarity between two images can equivalently be expressed in terms of dot products between their patch representations.
Let $\{\mathbf{f}_{i}^{A}\in\mathbb{R}^{512}\}_{i=1}^{K}$ and $\{\mathbf{f}_{i}^{B}\in\mathbb{R}^{512}\}_{i=1}^{K}$ denote the patch-level feature vectors for two face images, where $K$ is the number of patches and $\mathbf{f}_{i}^{A}$ is the feature of the $i$-th patch of an image $A$ and respectively $\mathbf{f}_{i}^{B}$ is the feature of the $i$-th patch of an image $B$.
The global representations are defined as the mean of the patch features:
@@
The similarity between the global representations is computed using cosine similarity:
@@
Expanding the numerator using (3):
$\mathbf{F}^{A}\cdot\mathbf{F}^{B}$ $=\left(\frac{1}{K}\sum_{i=1}^{K}\mathbf{f}_{i}^{A}\right)\cdot\left(\frac{1}{K}\sum_{j=1}^{K}\mathbf{f}_{j}^{B}\right)$ (5)
$=\frac{1}{K^{2}}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{A}\cdot\mathbf{f}_{j}^{B}.$
Similarly, the squared norms of the global representations are:
$\|\mathbf{F}^{A}\|^{2}$ $=\frac{1}{K^{2}}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{A}\cdot\mathbf{f}_{j}^{A},$ (6)
$\|\mathbf{F}^{B}\|^{2}$ $=\frac{1}{K^{2}}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{B}\cdot\mathbf{f}_{j}^{B}.$ (7)
Substituting these into (4) yields a similarity metric expressed in terms of dot products between patch representations:
@@ IV Experiment
Dataset. We evaluate the verification performance of the proposed method on seven benchmark datasets. LFW [58], CFP-FP [59], CPLFW [60], CALFW [61], and AGEDB [62] are commonly used in recent face recognition research and provide high-quality images for assessing robustness to age and pose variations. In addition, the Eclipse (ECL) and Hadrian (HAD) datasets [63] are used to evaluate performance under variations in illumination and facial hair. We follow the commonly utilized approach, that is, 10-fold cross-validation for all dataset to report verification rates.
TABLE II: Face verification accuracy (%) across seven benchmark datasets for each receptive field size and position individually. Additionally, score-level combination of receptive fields are given at the bottom. Results are reported for receptive fields of sizes $28\times 28$, $56\times 56$, and $112\times 112$, with coordinates $(x,y)$ indicating the top-left corner position. The best-performing receptive field for $56\times 56$ is highlighted in red, while the top-3 performing fields for $28\times 28$ are shown in blue for each dataset. Average accuracy rates across seven datasets are given in the last column.
@@
Comb.
$28\times 28$
- $99.67$ $96.99$ $90.32$ $96.03$ $95.20$ $80.05$ $93.60$ $93.12$
Comb.
$56\times 56$
- $99.82$ $99.10$ $93.70$ $97.82$ $95.97$ $84.20$ $95.65$ $95.18$
Comb.
$28\times 28$
$56\times 56$
- $99.73$ $99.17$ $93.65$ $98.03$ $96.03$ $85.12$ $96.87$ $95.51$
In our experiments, we use two receptive field sizes, $56\times 56$ and $28\times 28$, on $112\times 112$ face images to evaluate the verification performance. The layouts of these receptive fields are illustrated in Fig. 3. We first analyze the verification performance of individual patches by training the CNNs described in Section III-A. Each model is trained for 20 epochs on the WebFace4M dataset [64], using the ResNet100 architecture [57]. For the $56\times 56$ patches, the network retains the same number of parameters as in [15] with $112\times 112$ inputs, as we use a stride of 1 instead of 2 in the first ResNet block. On the other hand, for the $28\times 28$ patch size, the feature map reduces to $4\times 4$ in the last ResNet block, resulting in fewer parameters in the fully-connected layer compared to the baseline approach [15].
IV-A Verification Rates for Region-based Similarity.
Individual verification performance of the each receptive field is given in Tab. II. Note that, while there are forty-nine different positions for $28\times 28$ patches, shown in Fig. 3, we have twenty-eight CNNs as corresponding patches on the left and right trained with the same network due to face similarity. Similarly six networks are used for $56\times 56$ patches instead of nine. For example, in Fig. 3, patches at $(28,28)$ and $(56,28)$ trained with the same CNNs. As we apply horizontal flipping to augment representations during inference following [65], only left-side positions are depicted in the Tab. II. Colored text shows the best-performing patches.
While patches in the middle region of the face achieve the highest accuracies, the lower half of the face outperforms the upper half across most datasets. However, on Hadrian (HAD) [63], patches in the upper regions perform better than those in the lower regions due to significant beard and mustache variations between image pairs. Although smaller patch sizes lead to an expected drop in accuracy, surprisingly, a single $56\times 56$ patch at position $(28,28)$ achieves verification rates comparable to the full $112\times 112$ receptive field.
Combination of Receptive Fields. Score-level combination of the receptive fields is used for verification decisions. To compare the performance of the different receptive field sizes, combinations are performed in 3 ways: only $28\times 28$ patches, only $56\times 56$ and combination of $28\times 28$ and $56\times 56$. Verification rates using the full image size $112\times 112$ [15], with the same number of CNNs, are also given for comparison. Results are shown at the bottom of Tab II.
@@
RRFNet-28
(40% mask)
ResNet50 Casia-WebFace $99.33$ $97.71$ $90.37$ $95.18$ $93.77$ $72.28$ $81.73$ $90.05$
RRFNet-28
(20% mask)
ResNet50 Casia-WebFace $99.43$ $97.61$ $90.13$ $94.73$ $94.03$ $73.57$ $83.43$ $90.42$
RRFNet-28
(w/o mask)
ResNet50 Casia-WebFace $99.33$ $97.73$ $89.73$ $95.28$ $94.22$ $73.40$ $81.00$ $90.10$
RRFNet-56
(40% mask)
ResNet50 Casia-WebFace $99.48$ $98.11$ $90.28$ $95.70$ $94.25$ $73.92$ $85.73$ $91.07$
RRFNet-56
(20% mask)
ResNet50 Casia-WebFace $99.55$ $97.73$ $91.13$ $96.03$ $94.47$ $74.35$ $84.93$ $91.17$
RRFNet-56
(w/o mask)
ResNet50 Casia-WebFace $99.60$ $97.86$ $90.95$ $95.65$ $94.20$ $75.42$ $84.95$ $91.23$
@@
RRFNet-28
(40% mask)
ResNet100 WebFace4M $99.78$ $99.23$ $94.67$ $98.03$ $96.05$ $82.85$ $94.30$ $94.99$
RRFNet-28
(20% mask)
ResNet100 WebFace4M $99.73$ $99.19$ $94.75$ $97.97$ $95.88$ $83.70$ $95.17$ $95.20$
RRFNet-28
(w/o mask)
ResNet100 WebFace4M $99.77$ $99.30$ $94.57$ $98.02$ $96.02$ $83.37$ $94.47$ $95.07$
RRFNet-56
(40% mask)
ResNet100 WebFace4M $99.83$ $99.39$ $95.25$ $98.18$ $96.08$ $84.55$ $96.52$ $95.69$
RRFNet-56
(20% mask)
ResNet100 WebFace4M $99.82$ $99.36$ $95.10$ $98.18$ $95.97$ $84.58$ $96.13$ $95.59$
RRFNet-56
(w/o mask)
ResNet100 WebFace4M $99.75$ $99.37$ $94.98$ $98.03$ $95.97$ $84.28$ $95.80$ $95.46$

### v2

@@ Abstract
Index Terms:
Face Recognition, Biometrics, Computer Vision.
@@ I Introduction
As opposed to post-hoc approaches, interpretable models emphasize constrained design choices to make decisions inherently more interpretable [13, 26, 27, 28, 29, 30]. One prominent example is ProtoPNet [31], which represents model outputs as a weighted sum of similarities between the input and prototypes learned from the training images. This formulation enables localized and case-based reasoning by relating regions of the input image to the prototypes. However, its applicability is limited to closed-set recognition, as the prototypes are drawn from training data.
In this work, we introduce RRFNet to optimize face recognition through the fusion of patch-level similarity scores. Unlike the traditional approach of learning a single feature vector for a similarity assessment, RRFNet defines global similarity as the sum of local similarities (illustrated in Fig. 1). Since this formulation decomposes similarity into contributions from sub-regions, it yields saliency maps as an inherent part of the decision process, eliminating the need for computationally expensive post-hoc approaches.
We evaluate our approach using two patch sizes, $28\times 28$ and $56\times 56$, on $112\times 112$ face images. Although restricted receptive fields constrain the available spatial context, we surprisingly find that this design choice improves verification accuracy compared to state-of-the-art methods when $56\times 56$ patches are used for similarity measurement. Even with patches as small as $28\times 28$, our method achieves competitive verification performance. Note, performing verification based on local similarities inherently provides regional explanations, leading to more interpretable decision-making compared to global feature extraction approaches.
This paper is organized as follows. In Section II, we review the differences between post-hoc explanation methods and inherently interpretable approaches. Section III introduces two approaches for measuring similarity from patch representations. First, in Section III-A, we define a global distance metric formulated as a weighted sum of region-based similarities. Next, in Section III-B, we present RRFNet and show that with a small modification to the ResNet architecture, similarity decisions become more interpretable in terms of patch-level similarities. Experimental results for both methods are reported in Section IV, followed by quantitative analyses on explainability in Section V. Finally, Section VI presents conclusions and discusses directions for future research.
@@ II Related Work
Refer to caption (0,0) Refer to caption (14,0) Refer to caption (28,0) Refer to caption (42,0) Refer to caption (56,0) Refer to caption (70,0) Refer to caption (84,0) Refer to caption (0,14) Refer to caption (14,14) Refer to caption (28,14) Refer to caption (42,14) Refer to caption (56,14) Refer to caption (70,14) Refer to caption (84,14) Refer to caption (0,28) Refer to caption (14,28) Refer to caption (28,28) Refer to caption (42,28) Refer to caption (56,28) Refer to caption (70,28) Refer to caption (84,28) Refer to caption (0,42) Refer to caption (14,42) Refer to caption (28,42) Refer to caption (42,42) Refer to caption (56,42) Refer to caption (70,42) Refer to caption (84,42) Refer to caption (0,56) Refer to caption (14,56) Refer to caption (28,56) Refer to caption (42,56) Refer to caption (56,56) Refer to caption (70,56) Refer to caption (84,56) Refer to caption (0,70) Refer to caption (14,70) Refer to caption (28,70) Refer to caption (42,70) Refer to caption (56,70) Refer to caption (70,70) Refer to caption (84,70) Refer to caption (0,84) Refer to caption (14,84) Refer to caption (28,84) Refer to caption (42,84) Refer to caption (56,84) Refer to caption (70,84) Refer to caption (84,84) (a) $28\times 28$ Refer to caption (0,0) Refer to caption (28,0) Refer to caption (56,0) Refer to caption (0,28) Refer to caption (28,28) Refer to caption (56,28) Refer to caption (0,56) Refer to caption (28,56) Refer to caption (56,56) (b) $56\times 56$ Refer to caption (c) $112\times 112$
@@
A great amount of research has focused on post-hoc explanation approaches, where the decision of a black-box model is explained by generating heatmaps that highlight important pixels that influence the model’s predictions most. These visual explanations can help identify relevant regions, offering insights into where the model looks when making decisions. However, evaluating the quality of these explanations remains an unsolved challenge. Furthermore, post-hoc methods can be misleading, as similar heatmaps are often generated for both correct and incorrect predictions, undermining the reliability of these explanations [45, 46, 47, 48, 49, 9]. While it has been shown that searching sub-regions can increase faithfulness of the explanations [50, 51, 52], computational overhead limits the practical deployment.
@@
Another line of research in recent years focuses on incorporating interpretability directly into the vision models [53, 31, 54, 55, 56]. The main component of these models is to decompose images into smaller parts and make predictions based on a weighted combination of these parts. In [31], they employ a prototype learning approach to explain the decision through part-based evidence. Prototypes are assigned to each class based on representative image patches from the training set, and the final prediction is calculated as a weighted sum of similarity to these prototypes. While this part-based evidence framework enhances interpretability, it is restricted to classes present in the training set, limiting its generalization to unseen categories.
Recent works in face recognition have explored the use of post-hoc explanation approaches [21, 22, 23, 24, 57, 58]. In contrast, our work aims to make decisions inherently interpretable by leveraging local similarities. Similar to our approach, a self-interpretable face recognition model in [56] is trained to provide patch-level similarity scores alongside its predictions. They introduce an xCos module to highlight the model’s regional attention, while the network’s receptive field still covers the entire image. In contrast, our method explicitly restricts the receptive field, ensuring that network decisions are directly determined by patch-level similarities. Furthermore, we quantitatively demonstrate that our approach performs significantly better on faithfulness metrics (see Section V).
Patch-level Feature Learning. While global feature learning approaches have shown great success in recent years [59, 60], BagNet [44] investigates the use of patch representations for object recognition. It restricts the receptive field of a CNN to enable local feature learning, while it results in a significant drop in accuracy. In face recognition, patch-level representations were extensively studied prior to the deep learning era [61, 62, 63, 64, 65, 66, 67, 68]. The typical approach involves extracting local features using hand-crafted methods and then fusing information from different regions. For instance, [67] combines the outputs of nearest neighbor classifiers trained on different face patches. However, it uses the same set of subjects for both training and testing, lacking an evaluation on unseen subjects. Similarly, while [68] uses hierarchical patch relationships to enhance local matching, its performance evaluation remains limited to fixed gallery subjects. In contrast to the traditional local feature extraction strategies, recent methods have shifted toward global feature learning [15, 14, 69, 70, 16] to achieve state-of-the-art verification performance. In this work, we demonstrate that feature learning does not need to operate at a global level to achieve high accuracy rates. Unlike traditional fusion methods, RRFNet integrates local feature learning with score fusion, demonstrating performance competitive with state-of-the-art methods.
@@ III Methodology
In this work, we propose a distance metric to decompose face similarity measure into local regions. For an image size of $W\times H$, we choose $w<W$ and $h<H$ to learn local representations for each $w\times h$ restricted patch. The image patches are uniformly distributed across the image. Two approaches are considered for learning local representations and constructing global image similarity metric: (i) training region-based CNNs for each $k$ patches (see Section III-A) and (ii) single CNN to learn a global image representation as the mean of local representations (see Section III-B). In the first approach, we adopt the same training objective as in [15] to learn local features for each patch obtained from $W\times H$ face images. After the CNNs are trained, a second training phase learns weights for each region similarity using logistic regression. Weighted sum of $k$ local similarity scores, obtained from corresponding regions between two images, is used to the build global distance metric. In the second approach, rather than training separate CNNs, a single CNN is trained and the global representation is computed as the mean of the patch-level features, enabling face similarity to be measured directly from patch representations. We use the terms patch and restricted receptive field interchangeably throughout the manuscript to describe our approach.
III-A Patch Representation Learning
We first divide each face image into a set of smaller, uniformly distributed patches. Each patch is treated as an independent unit, allowing the model to capture fine-grained localized feature representations. Unlike the traditional approach of generating a single $N$-dimensional representation for an entire image, we obtain $(k,N)$-dimensional representation where a face image is divided into $k$ patches. Patches are defined with position and window size. For a given face image, we extract $k$ patches, which we will refer to as $P_{1},P_{2},\dots,P_{k}$, each corresponding to a distinct region and size of the image. While $112\times 112$ is traditionally used image size for face recognition research in recent works, we analyze two patch sizes, $28\times 28$ and $56\times 56$ in our experiments, demonstrated in Fig. 2.
@@ III-A Patch Representation Learning
Let the output of each CNN, trained in the first step, corresponding to the $i$-th patch be $f_{i}$, which is a $N$-dimensional feature vector representing patch $P_{i}$. Local similarity between two images is computed for each corresponding patches $P_{i}^{A}$ from image $A$ and $P_{i}^{B}$ image $B$ using cosine similarity:
@@ III-B Restricted Receptive Field Network (RRFNet)
Comparison of the proposed RRFNet approach with the ResNet [60] is depicted in Table I. While traditional approach with ResNet produce a $512$-dimensional representation from the entire face image, our approach extracts $512$-dimensional representations for each restricted receptive field. At the end of the RRFNet, a global representation is obtained as the mean of the patch representations enabling expression of the global similarity as patch similarities (illustrated in Fig. 3).
Patch-level Similarity Metric. Typically, the cosine similarity between two $512$-dimensional face representations, extracted from a pretrained recognition model, is used for face verification. In our approach, the global face representation is defined as the mean of the patch-level representations. This formulation implies that the similarity between two images can equivalently be expressed in terms of the sum of dot products between all pairs of their patch representations.
Let $\{\mathbf{f}_{i}^{A}\in\mathbb{R}^{512}\}_{i=1}^{K}$ and $\{\mathbf{f}_{i}^{B}\in\mathbb{R}^{512}\}_{i=1}^{K}$ denote the patch-level feature vectors extracted from two face images, where $K$ is the total number of patches. Here, $\mathbf{f}_{i}^{A}$ and $\mathbf{f}_{i}^{B}$ represent the feature of the $i$-th patch in images $A$ and $B$, respectively.
The global representation of each image is obtained by averaging its patch features:
@@
The cosine similarity between the two global representations is defined as
@@
Substituting (3) into the numerator of (4), we obtain
$\mathbf{F}^{A}\cdot\mathbf{F}^{B}$ $=\left(\frac{1}{K}\sum_{i=1}^{K}\mathbf{f}_{i}^{A}\right)\cdot\left(\frac{1}{K}\sum_{j=1}^{K}\mathbf{f}_{j}^{B}\right)$
$=\frac{1}{K^{2}}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{A}\cdot\mathbf{f}_{j}^{B}.$ (5)
And the norms of the global representations can be written as
$\|\mathbf{F}^{A}\|$ $=\frac{1}{K}\sqrt{\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{A}\cdot\mathbf{f}_{j}^{A}},$ (6)
$\|\mathbf{F}^{B}\|$ $=\frac{1}{K}\sqrt{\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbf{f}_{i}^{B}\cdot\mathbf{f}_{j}^{B}}.$ (7)
Substituting (III-B)–(7) into (4), the scalar factors of $\tfrac{1}{K^{2}}$ cancel out, yielding a similarity formulation expressed entirely in terms of patch-level dot products:
@@ IV Experiment
Dataset. We evaluate the verification performance of the proposed method on seven benchmark datasets. LFW [71], CFP-FP [72], CPLFW [73], CALFW [74], and AGEDB [75] are commonly used in recent face recognition research for assessing robustness to age and pose variations. In addition, the Eclipse (ECL) and Hadrian (HAD) datasets [76] are used to evaluate performance under variations in illumination and facial hair.
In our experiments, we use two receptive field sizes, $56\times 56$ and $28\times 28$, on $112\times 112$ face images to evaluate the verification performance. We first analyze the verification performance of individual patches by training the CNNs described in Section III-A. Each model is trained for 20 epochs on the WebFace4M dataset [77], using the ResNet100 architecture [60]. For the $56\times 56$ patches, the network retains the same number of parameters as in [15] with $112\times 112$ inputs, as we use a stride of 1 instead of 2 in the first ResNet block. On the other hand, for the $28\times 28$ patch size, the feature map reduces to $4\times 4$ in the last ResNet block, resulting in fewer parameters in the fully-connected layer compared to the baseline approach [15].
IV-A Verification Rates for Region-based Similarity.
Individual verification performance of the each receptive field is given in Table II. Note that, while there are forty-nine different positions for $28\times 28$ patches, shown in Fig. 2, twenty-eight CNNs are trained because symmetric patches on the left and right sides share the same network due to facial symmetry. Similarly six networks are used for $56\times 56$ patches instead of nine. For example, in Fig. 2, patches at $(28,28)$ and $(56,28)$ trained with the same CNNs. As we apply horizontal flipping to augment representations during inference following [78], only left-side positions are depicted in the Table II.
While patches in the middle region of the face achieve the highest accuracies, the lower half of the face outperforms the upper half across most datasets. However, on Hadrian, patches in the upper regions perform better than those in the lower regions due to significant beard and mustache variations between image pairs. Although smaller patch sizes lead to an expected drop in accuracy, surprisingly, a single $56\times 56$ patch at position $(28,28)$ achieves verification rates comparable to the full $112\times 112$ receptive field.
Combination of Receptive Fields. Score-level combination of the receptive fields is used for verification decisions. To compare the performance of the different receptive field sizes, combinations are performed in 3 ways: only $28\times 28$ patches, only $56\times 56$ and combination of $28\times 28$ and $56\times 56$. Verification rates using the full image size $112\times 112$ [15], are also given for comparison. Results are shown at the bottom of Tab II.
TABLE II: Face verification accuracy (%) across seven benchmark datasets for each receptive field size and position individually. Additionally, score-level combination of receptive fields are given at the bottom. Results are reported for receptive fields of sizes $28\times 28$, $56\times 56$, and $112\times 112$, with coordinates $(x,y)$ indicating the top-left corner position. The best-performing receptive field for $56\times 56$ and the top-3 performing receptive fields for $28\times 28$ are highlighted in bold for each dataset.
@@
Comb. $28\times 28$ - $99.67$ $96.99$ $90.32$ $96.03$ $95.20$ $80.05$ $93.60$ $93.12$
Comb. $56\times 56$ - $99.82$ $99.10$ $93.70$ $97.82$ $95.97$ $84.20$ $95.65$ $95.18$
Comb. $28\times 28$ & $56\times 56$ - $99.73$ $99.17$ $93.65$ $98.03$ $96.03$ $85.12$ $96.87$ $95.51$
@@
BagNet-17 [44] ResNet50 Casia-WebFace $98.35$ $92.53$ $83.57$ $89.32$ $90.18$ $60.87$ $67.82$ $83.23$
BagNet-33 [44] ResNet50 Casia-WebFace $98.98$ $94.93$ $87.33$ $92.63$ $92.68$ $66.10$ $72.27$ $86.42$
RRFNet-28 (40% mask) ResNet50 Casia-WebFace $99.33$ $97.71$ $90.37$ $95.18$ $93.77$ $72.28$ $81.73$ $90.05$
RRFNet-28 (20% mask) ResNet50 Casia-WebFace $99.43$ $97.61$ $90.13$ $94.73$ $94.03$ $73.57$ $83.43$ $90.42$
RRFNet-28 (w/o mask) ResNet50 Casia-WebFace $99.33$ $97.73$ $89.73$ $95.28$ $94.22$ $73.40$ $81.00$ $90.10$
RRFNet-56 (40% mask) ResNet50 Casia-WebFace $99.48$ $98.11$ $90.28$ $95.70$ $94.25$ $73.92$ $85.73$ $91.07$
RRFNet-56 (20% mask) ResNet50 Casia-WebFace $99.55$ $97.73$ $91.13$ $96.03$ $94.47$ $74.35$ $84.93$ $91.17$
RRFNet-56 (w/o mask) ResNet50 Casia-WebFace $99.60$ $97.86$ $90.95$ $95.65$ $94.20$ $75.42$ $84.95$ $91.23$
@@
BagNet-17 [44] ResNet100 WebFace4M $99.63$ $97.89$ $91.47$ $95.50$ $94.98$ $75.33$ $82.63$ $91.06$
BagNet-33 [44] ResNet100 WebFace4M $99.67$ $98.81$ $93.28$ $96.78$ $95.62$ $78.92$ $86.15$ $92.75$
RRFNet-28 (40% mask) ResNet100 WebFace4M $99.78$ $99.23$ $94.67$ $98.03$ $96.05$ $82.85$ $94.30$ $94.99$
RRFNet-28 (20% mask) ResNet100 WebFace4M $99.73$ $99.19$ $94.75$ $97.97$ $95.88$ $83.70$ $95.17$ $95.20$
RRFNet-28 (w/o mask) ResNet100 WebFace4M $99.77$ $99.30$ $94.57$ $98.02$ $96.02$ $83.37$ $94.47$ $95.07$
RRFNet-56 (40% mask) ResNet100 WebFace4M $99.83$ $99.39$ $95.25$ $98.18$ $96.08$ $84.55$ $96.52$ $95.69$
RRFNet-56 (20% mask) ResNet100 WebFace4M $99.82$ $99.36$ $95.10$ $98.18$ $95.97$ $84.58$ $96.13$ $95.59$
RRFNet-56 (w/o mask) ResNet100 WebFace4M $99.75$ $99.37$ $94.98$ $98.03$ $95.97$ $84.28$ $95.80$ $95.46$
TABLE IV: Percentage of patches in genuine pairs whose most similar match occurs at the same spatial location.
Dataset LFW CFPFP CPLFW AGEDB CALFW ECL HAD
Same region (%) $46.2$ $20.3$ $23.6$ $34.6$ $39.9$ $48.9$ $49.6$